If (9, -2) is the endpoint of a line segment, and (6, 2) is its midpoint, find the other endpoint.
a. (3, 6)
b. (17, -8)
c. (3, -6)
d. (15, -10)
I spend an hour trying ti figure it out ( please help
thanks
Michael

Answer by Cromlix(4381)

Hi there,
If you give the line points.
A (you are trying to find)
B (Mid pt) (6,2)
C (End pt) (9,-2)
AB/BC = b - a/c - b = 1/1
Cross multiply
b - a = c - b
-a = c - b - b
-a = c -2b
Multiply both sides by -1
a = 2b - c
a = 2(6,2) - (9,-2)
a = (12,4) - (9,-2)
a = (3, 6)
Therefore point A = (3,6)
Hope this helps:-)